OpenStudy (anonymous):

I'm having a hard time figuring out what to do with the -2, since the denominators are equivalent already. What is my next step? Thanks! 7 2-3x ---- -2 = ----- x+5 x+5

OpenStudy (anonymous):

Common denom. x+5 -> 2 ( x+5 )

OpenStudy (anonymous):

\[\frac{ 7 }{ x+5} -2 = \frac{ 2-3x }{ x+5 }\] So your saying: \[(2)\frac{ 7 }{ x+5} = (2)\frac{ 2-3x }{ x+5 }\] \[\frac{ 14 }{ 2(x+5)} = \frac{ 4-6x }{ 2(x+5) }\]? if so, why 2 and not -2?

OpenStudy (anonymous):

Which part you multiply 2 ( x +5) = ....?

OpenStudy (anonymous):

-> 7 - 2x - 10 = 2 - 3x - 2x - 3 = 2 + 3x 3x - 2x = 2 + 3 x = 5